NASA Space Station On-Orbit Status 3 March 2017

©NASA

NASA Space Station On-Orbit Status 3 March 2017.

Three Expedition 50 crew members recently surpassed 100 days in space and will come home in June. Also, SpaceX Dragon external experiments are being unloaded for installation on the International Space Station.

NASA astronaut Peggy Whitson has been living in space with her Expedition 50 crewmates for over 100 days now. She and Thomas Pesquet from the European Space Agency and Oleg Novitskiy from Roscosmos will return to Earth June 2 after 197 days in space. The trio launched Nov. 17 beginning a two-day trip to the station.

Whitson has reached out to her supporters on the ground and mentioned them with her "NASA Village" campaign from space. She launched to the station with posters signed by hundreds of support personnel and recently showed them off on orbit.

Watch Peggy Whitson describe the "NASA Village"
Visit Peggy Whitson's Facebook page

A pair of Earth observation experiments delivered last week on the SpaceX Dragon have been robotically removed from the resupply ship. Both experiments will be installed on specialized pallets on the outside of the orbital laboratory and activated for years of research.

The first experiment, SAGE-III, will observe how tiny particles interact with Earth's sunscreen, or ozone, possibly affecting the climate. The second, Lightning Image Sensor, will monitor lightning around the globe to improve weather forecasting, enhance climate models and increase aviation safety.

On-Orbit Status Report

Fluid Shifts Dilution Measurements: The 48S subjects initiated their Return minus 45 (R-45) Fluid Shifts Dilution Measurements activities, beginning at wakeup the subject collected saliva, blood and urine and insert them into Minus Eighty Degree Celsius Laboratory Freezer for ISS (MELFI) prior to ingesting a Sodium Bromide (NaBr) tracer. Throughout the day, the crew performed more urine, blood, and saliva collections, inserting those samples into MELFI as well. The crew configured for tomorrow's Baseline Imaging in the USOS. Fluid Shifts is divided into three experiment portions - Dilution Measurements, Baseline Imaging, and Baseline Imaging with Chibis (Lower Body Negative Pressure). The Fluid Shifts experiment investigates the causes for severe and lasting physical changes to astronaut's eyes. Because the headward fluid shift is a hypothesized contributor to these changes, reversing this fluid shift with a lower body negative pressure device is investigated as a possible intervention. Results from this study may help to develop preventative measures against lasting changes in vision and eye damage.

Energy Requirements for Long-Term Space Flight (Energy): The crew performed Day 2 of the 11-day Energy protocol, continuing with urine collection and stowage of deployed Pulmonary Function System (PFS) equipment. Throughout the day, the crew continued to log their food and drink consumptions as well. The Energy investigation measures adaptations in the components of total energy expenditure of crewmembers, so that an equation to determine the energy requirements of crewmembers during space flight may be derived. Such knowledge of energy requirements is of a great importance to ensure health, good performance and the overall success of a mission, and also contributes to ensure adequate exercise load and cargo allotments for food during space flight.

Advanced Plant Experiments (APEX)-04 Photo: The crew retrieved an APEX-04 petri plate and photographed it on the Advanced Biology Research Facility (ABRS) photogrid. The picture will be downlinked for ground teams to evaluate the growth status. The APEX-04 (Epigenetic change in Arabidopsis thaliana in response to spaceflight - differential cytosine DNA methylation of plants on the ISS) experiment investigates how spaceflight can affect an organism on a molecular level without altering its outward appearance. Previous research has demonstrated that large numbers of plant genes are expressed differently in space compared to plants on Earth, which can alter how plants grow. This investigation studies the entire genome of thale cress plants grown in space, creating maps of spaceflight-specific changes in certain groups of genes. Results give new insight into plants' molecular responses to spaceflight, which benefits efforts to grow plants in space for food and oxygen.

Pressurized Mating Adapter 3 (PMA3) Ingress: The crew ingressed PMA3. They removed the Common Berthing Mechanism (CBM) Center Disk Cover to provide access to the vestibule areas. They then ingressed PMA3 and removed stowage which needs to be removed prior to the PMA3 relocation. The crew also scavenged the Node 3 Port Forward Intermodular Ventilation (IMV) Valve. There are no spare IMV Valves on orbit currently. Due to the PMA3 ingress and stowage relocation out of PMA3, the removal of this valve became much more feasible for access and crew time requirements. When the crew is complete with today's PMA3 activities they will close the Node 3 Port Hatch. Over the next few weeks the crew will continue to prepare the PMA-3 to be relocated from Node 3 Port to Node 2 Zenith for its new job as an IDA docking port for ISS.

Mobile Servicing System (MSS) Operations: Overnight, the Robotics Ground Controllers maneuvered the Space Station Remote Manipulator System (SSRMS) and Special Purpose Dexterous Manipulator (SPDM) to extract the Stratospheric Aerosol and Gas Experiment (SAGE) Instrument Panel (IP) from the SpX-10 Dragon Trunk. They then maneuvered the SPDM and SSRMS to install the SAGE IP on the SPDM Enhanced Orbital Replacement Unit (ORU) Temporary Platform (EOTP). Finally the Robotics Ground Controllers rotated the SPDM EOTP and reconfigured the SPDM for the extraction of the SAGE Nadir Viewing Platform (NVP) from the Dragon Trunk. This afternoon, they extracted the Stratospheric Aerosol and Gas Experiment (SAGE) Nadir Viewing Platform (NVP) from the SpX-10 Dragon Trunk. They then maneuvered the SPDM and SSRMS to remove the Optical Payload for Lasercomm Science (OPALS) from the SPDM EOTP. Later today, SSRMS & SPDM will install OPALS into the Dragon trunk. MSS performance was nominal..

Today's Planned Activities
All activities were completed unless otherwise noted.

Fluid Shifts Saliva Baseline Collection
FLUID SHIFTS. Urine Sample Collection
FLUID SHIFTS. Sample MELFI Insertion
FLUID SHIFTS. Basic blood collection
ENERGY Diet Log of Breakfast
FLUID SHIFTS. Galley Water Collection and Tracer Ingestion Ops
Fluid Shifts Refrigerated Centrifuge Configuration
FLUID SHIFTS. Blood sample removal from refrigerated centrifuge
Replacement of БД-2 Treadmill Belt.
Energy PFS Conclude
MSG Clean Up
Waste Hygiene Compartment (WHC) Piping R&R
Regeneration of БМП Ф2 Micropurification Cartridge (start)
Node 3 Endcone Cleanout
Photo TV Pressurized Mating Adapter 3 (PMA 3) Video Setup
PMA3 Ingress
PMA 3 CPA Locate
APEX-04 Petri Plate Photo
FLUID SHIFTS. 3-hour blood sample collection
Fluid Shifts Blood Collection - Operator
Structures and Mechanisms Center Disk Cover Removal
Fluid Shifts Refrigerated Centrifuge Configuration
FLUID SHIFTS. 3-hour saliva sample collection
Pressurized Mating Adapter 3 (PMA 3) Stowage Removal
FLUID SHIFTS. Sample MELFI Insertion
Fine Motor Skills iPad App Update
Habitability Human Factors Directed Observations - Subject
Connecting HDV Sony HVR-Z7E Camcorder in SM and Running NASA MPEG-2 Viewer
ENERGY Water Sample
Energy Urine Sample Collection
IFM Node 3 Port Intermodule Ventilation Cap Installation
FLUID SHIFTS. 5-hour Saliva Sample
Node 3 Port Forward IMV Valve Scavenge
Energy Urine Sample Collection
Fluid Shifts CCFP Baseline Setup
Pressurized Mating Adapter 3 (PMA3) Node 3 Port Hatch Close
Multi-purpose Small Payload Rack (MSPR) /Group Combustion Module (GCM) Component Deactivation
Battery Stowage Assembly (BSA) Operation Termination
Replacement of БД-2 Treadmill Belt.
Battery Stowage Assembly (BSA) Operation Initiation
ENERGY Diet Log of Lunch
PILOT-T. Preparation for the experiment.
Fluid Shifts Urine Collection Stow
Deactivation of HDV Sony HVR-Z7E camcorder and closing of NASA MPEG-2 Viewer application
Public Affairs Office (PAO) High Definition (HD) Config LAB Setup
Strata Status Check
PILOT-T. Experiment Ops.
Long Duration Sorbent Testbed Inlet Filter Change.
Public Affairs Office (PAO) Event in High Definition (HD) - Lab
FLUID SHIFTS. HW batteries installation
Long Duration Sorbent Testbed Status Check.
Polar Hardware Install into Express Rack
Polar Desiccant Pack Swap
Wastewater Storage Tank Assembly (WSTA) Fill
PILOT-T. Closeout Ops.
СОЖ maintenance
Space Headaches - Weekly Questionnaire
NanoRacks Module-40 Ops in Cupola
Microgravity Expanded Stem Cells Wash Kit MERLIN Insertion
URAGAN. Observation and photography using Photo Equipment
Regeneration Micropurification Unit (БМП) Ф2 Absorption Cartridge (end)
ENERGY Diet Log of Dinner

Completed Task List Items
EVA H2O Hardware Label
PIP COL SUP1 Hardware Deploy
CDM Data Download

Ground Activities
All activities were completed unless otherwise noted.
SAGE Extraction
OPALS Transfer

Three-Day Look Ahead:
Saturday, 03/04: Weekly Housekeeping, Microgravity Expanded Stem Cells (MESC) Session
Sunday, 03/05: Microgravity Expanded Stem Cells (MESC) Session
Monday, 03/06: EVA Tool Config, Fluid Shifts, MESC Hardware Teardown

QUICK ISS Status - Environmental Control Group:
Component - Status
Elektron - On
Vozdukh - Manual
[СКВ] 1 - SM Air Conditioner System ("SKV1") - On
[СКВ] 2 - SM Air Conditioner System ("SKV2") - Off
Carbon Dioxide Removal Assembly (CDRA) Lab - Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 - Operate
Major Constituent Analyzer (MCA) Lab - Operate
Major Constituent Analyzer (MCA) Node 3 - Operate
Oxygen Generation Assembly (OGA) - Process
Urine Processing Assembly (UPA) - Process
Trace Contaminant Control System (TCCS) Lab - Off
Trace Contaminant Control System (TCCS) Node 3 - Full Up

Please follow SpaceRef on Twitter and Like us on Facebook.